Constants A cylindrical bucket, open at the top, is 25.0 cm high and 15.0 cm in diameter. A circular hole with a cross-sectional area 1.99 cm2 is cut in the center of the bottom of the bucket. Water flows into the bucket from a tube above it at the rate of 2.25x10 4 m3/s
Answer #1

Solution)

Let height of water be H m

We know, speed

v=sqrt(2gh)

Also, we can conclude

v*1.99e-4=2.5e-4

so sqrt(2gh)*1.43e-4=2.25e-4

h=0.126 m (Ans)

==========

Good luck!:)